### v1

# Minimum Weighted Feedback Arc Sets for Ranking from Pairwise Comparisons

Abstract
The Minimum Weighted Feedback Arc Set (MWFAS) problem is fundamentally connected to the Ranking Problem - the task of deriving global rankings from pairwise comparisons. Recent work [He et al. ICML2022] has advanced the state-of-the-art for the Ranking Problem using learning-based methods, improving upon multiple previous approaches. However, the connection to MWFAS remains underexplored. This paper investigates this relationship and presents efficient combinatorial algorithms for solving MWFAS, thus addressing the Ranking Problem. Our experimental results demonstrate that these simple, learning-free algorithms not only significantly outperform learning-based methods in terms of speed but also generally achieve superior ranking accuracy. [^1] This paper is a preliminary report.
1 Introduction
The ranking problem and the Minimum Weighted Feedback Arc Set (MWFAS) problem are closely related; however, the relationship between them has not been extensively explored in the literature. Our review of existing studies on ranking shows that only a few papers, such as [14] and [7], briefly mention the feedback arc set problem and a polynomial-time approximation scheme (PTAS) for tournaments [21], without conducting a thorough analysis of the connection between ranking and MWFAS. To better understand this relationship, we present definitions for both problems:
Definition: The Ranking Problem Input: A directed graph $G=(V,E)$, where $V$ is the set of nodes representing agents or items to be ranked, and $E$ is the set of directed edges. Each edge $(i,j)\in E$ has a weight $w_{ij}>0$, representing the strength of preference or the victory of node $i$ over node $j$ in a pairwise comparison. Output: A ranking $R:V\to\mathbb{N}$ that assigns a rank $R(i)$ to each node $i\in V$. The ranking should minimize the total inconsistency with the pairwise comparisons represented by $E$. Specifically, the goal is to minimize the number of violated edges (or minimize the total weight of violated edges) where $R(i)>R(j)$ despite an edge $(i,j)$ indicating a preference for $i$ over $j$.
Definition: The Minimum Weighted Feedback Arc Set Problem Input: A directed graph $G=(V,E)$, where $V$ is the set of nodes, and $E$ is the set of directed edges. Each edge $(i,j)\in E$ has an associated weight $w_{ij}>0$, representing the cost of removing that edge. Output: A subset of edges $E^{\prime}\subseteq E$ such that removing all edges in $E^{\prime}$ results in a Directed Acyclic Graph (DAG). The goal is to minimize the total weight of the removed edges, i.e., $\sum_{(i,j)\in E^{\prime}}w_{ij}$, thereby ensuring that the graph becomes acyclic with the least possible cost.
The relationship between the ranking problem and the MWFAS problem is rooted in the goal of resolving inconsistencies within a directed graph to derive a meaningful ordering of nodes. In the ranking problem, the objective is to generate a ranking that aligns as closely as possible with given pairwise comparisons, minimizing the inconsistencies where ranks contradict preference directions. In a similar vein, the MWFAS problem focuses on eliminating cycles in the graph by removing the minimum weight of edges, ultimately transforming the graph into a Directed Acyclic Graph (DAG).
This acyclic structure can then be used to perform a topological sort, yielding a ranking that adheres to the remaining pairwise comparisons. Therefore, the ranking problem can be reformulated as an instance of the MWFAS problem, wherein eliminating cycles allows for a consistent ranking that reflects the original preferences as much as possible. It is worth noting that the Minimum Feedback Arc Set problem, even in its unweighted version, is NP-hard [20], highlighting the inherent computational difficulty in deriving consistent rankings.
1.1 Ranking Problem and Minimum Weighted Feedback Arc Set (MWFAS)
The ranking problem can be effectively modeled as an instance of the Minimum Weighted Feedback Arc Set (MWFAS) problem. An edge from vertex $i$ to vertex $j$ with weight $w$ signifies that agent $i$ is preferred over agent $j$ by $w$ units. When the comparison graph is already a Directed Acyclic Graph (DAG), we can simply apply a topological sort to generate the ranking, where each edge from vertex $i$ to vertex $j$ ensures that $i$ is ranked higher than $j$.
In practice, however, comparison graphs often contain cycles, leading to inconsistencies in the pairwise preferences. To address these inconsistencies, we aim to delete the smallest possible subset of edges (in terms of total weight) such that the remaining graph is acyclic. This transformation corresponds directly to solving the MWFAS problem, where the goal is to remove edges with minimal cost to obtain an acyclic graph, thereby enabling a consistent ranking.
The MWFAS problem, like its unweighted counterpart, is known to be computationally challenging (NP-hard). While the Minimum Feedback Arc Set (MFAS) problem has been extensively studied in its unweighted form[29][22][12], the weighted version has received relatively less attention in the literature. The following sections discuss notable works aimed at solving the ranking problem.
1.2 Related Works
1.2.1 Previous Work About Minimum Weighted Feedback Arc Set
  • The Minimum Weighted Feedback Arc Set (MWFAS) problem has received significant attention, particularly regarding approximation algorithms. Early work by [10] provided an approximation algorithm for the MWFAS problem, followed by improvements by [9], that introduced a divide-and-conquer approach using spreading metrics. Although effective, these algorithms used the ellipsoid method for solving linear programming relaxations, making them computationally slow in practice.
    Later, [8] proposed an $O(VE)$ heuristic based on the local ratio method [3]. This heuristic was shown to be efficient in practice, providing a $\lambda$-approximation, where $\lambda$ is the length of the longest cycle in the graph.
    [2] presented an exact method for solving the Minimum Feedback Arc Set (MFAS) problem with a focus on sparse directed graphs. They proposed a lazy cycle enumeration approach to iteratively extend an incomplete cycle matrix until an optimal solution was found. Given the NP-hard nature of the problem, their exact algorithm has exponential runtime and does not perform efficiently for large graphs.
    [19] proposed an $O(VE^{4})$ algorithm for MWFAS. However, as the number of edges in the graph increases, the algorithm’s performance becomes slower. Our work implemented some of these algorithms, and the results, available in our GitHub repository, confirm the practical limitations regarding their speed.
    [26] addresses the parameterized complexity of the feedback vertex and arc set problems in tournaments, focusing on both weighted and unweighted versions. It presents fixed parameter tractable (FPT) algorithms for these problems, including an $O((2.4143)^{k}n^{\omega})$ algorithm for the weighted feedback vertex set problem and an $O((2.415)^{k}n^{\omega})$ algorithm for the weighted feedback arc set problem where $\omega$ is exponent of the running time of the best matrix multiplication algorithm, and $k$ is the size of the feedback set. The paper also introduces algorithms for the parametric duals of these problems and explores their complexity in oriented directed graphs and general directed graphs.
    [17] [28] explores the use of satisfiability-modulo theories (SMT) solvers to solve the Minimum Feedback Arc Set (MFAS) problem. It presents an SMT encoding for MFAS and highlights the use of incremental solving features, such as unsatisfiable cores. The results show that while the SMT-based approach has potential, it remains computationally infeasible for larger instances, with solving times that scale poorly. The authors also provide complexity bounds, demonstrating that the MFAS problem remains challenging even with advanced SMT techniques.
1.2.2 Previous Work About Ranking Problem
The ranking problem has also been studied extensively in the literature. [7] introduced Sync-Rank, a method that employs eigenvector and semidefinite programming synchronization to solve ranking, constrained ranking, and rank aggregation problems. The approach leverages spectral or SDP relaxations, followed by rounding procedures to recover rankings. Sync-Rank outperformed other algorithms in both synthetic and real-world datasets while supporting semi-supervised ranking and identifying consistent partial rankings. [6] addressed the problem of ranking players using incomplete pairwise comparison data within the Bradley-Terry-Luce model framework. It established the minimax rate for this ranking task with respect to Kendall’s tau distance, identifying a unique transition between exponential and polynomial rates based on the signal-to-noise ratio. The paper proposed a divide-and-conquer algorithm that achieved the minimax rate, demonstrating its theoretical optimality and computational efficiency.
[27] proposed a simple, robust, and optimal ranking algorithm based on the Copeland counting method for pairwise comparisons. The algorithm ranks items by counting pairwise comparison wins, which is computationally efficient and achieves substantial speed-ups compared to previous methods. Unlike other models such as the Bradley-Terry-Luce model, this method does not rely on parametric assumptions and is robust to different data types.
[25] presented the Distinguishing Feature (DF) model, which effectively learns from pairwise comparisons, even in cases involving intransitive preferences. The DF-Learn algorithm, which is based on a Siamese neural network, outperformed standard models on synthetic and real-world datasets while capturing complex cyclic dependencies.
[16] proposed a GNN-based pairwise surrogate model to accelerate evolutionary optimization by predicting the superiority of candidate solutions rather than evaluating their exact fitness values. This model significantly reduced the number of fitness evaluations while maintaining comparable optimization performance. It is designed to work with any comparison-based evolutionary algorithm, offering a computationally efficient alternative for handling expensive optimization problems.
[11] introduced Bias-Aware Ranker from Pairwise Comparisons (BARP), which extends the Bradley-Terry model with bias parameters to correct evaluator biases and produce unbiased rankings. BARP outperformed baseline methods on synthetic and real-world datasets.
[15] introduces a novel pipeline called Graph Neural Re-Ranking (GNRR), which integrates Graph Neural Networks (GNNs) into the re-ranking phase of information retrieval systems. By constructing corpus graphs to model semantic and lexical document relationships, the GNRR approach captures cross-document interactions, improving ranking performance on popular metrics compared to traditional uni-variate baselines. Experimental results show consistent performance gains, especially in challenging datasets, highlighting the potential of GNNs to refine document rankings effectively. The most similar work to ours is [5], which focuses on vertex ordering in directed graph streams. They address the Feedback Arc Set problem and propose removing feedback arcs for conducting topological sorts. However, unlike our work, their focus is on unweighted tournament graphs without experimental validation, whereas we address weighted general graphs with extensive experiments.
[4] presents exact algorithms for solving several vertex ordering problems on graphs. The proposed algorithms operate in either $O^{*}(2^{n})$ time and space or $O^{*}(4^{n})$ time and polynomial space, leveraging dynamic programming and recursive techniques similar to the Held-Karp and Gurevich-Shelah methods. These algorithms generalize approaches for the Travelling Salesman Problem and can be applied to a variety of vertex ordering problems, including treewidth and pathwidth. The main contribution lies in providing algorithms with moderately exponential complexity for a broad class of graph ordering problems.
2 A novel framework for ranking problem based on pairwise comparisons using MWFAS
Based on our analysis of the existing algorithms, we chose to implement the heuristic algorithm from [8] for our ranking algorithm due to its efficiency in practice. The local ratio-based heuristic provides an effective trade-off between solution quality and runtime performance, particularly for graphs with a large number of vertices and edges.
2.1 Highlights of Our Work
  • Speed: Our method is very fast. As shown in the results, our method provides an answer in seconds. [18] provided many tables; however, none of them showed the running time of the experiment. It was only claimed that if an experiment did not answer after one week, its performance in the table was filled with ”NaN”. [13] and [23] also did not provide the running time of their programs. To the best of our knowledge, our method is the fastest in this regard, and its performance is comparable to that of other methods.
  • Flexibility: Our method uses an algorithm to solve the MWFAS problem as a black box. The user can choose any algorithm for solving MWFAS, thus allowing them to handle the trade-off between running time and performance. If performance is more important, more precise algorithms can be used; otherwise, faster heuristics can be chosen.
  • Simplicity: Our method is very simple, using only well-known algorithms like topological sort, depth-first search (or breadth-first search) to find cycles, and heap data structures.
2.2 Algorithm
Input: Pairwise comparisons of elements as a weighted directed graph file.
Output: Rankings of elements.
Step 1: Read Graph and Initialize Structures;
Parse the input file to read edges $\{(u,v,w)\}$ representing directed edges with weights $w$;
Initialize an adjacency list representation for the graph;
Store edge weights in a dictionary;
Step 2: Find Minimum Weighted Feedback Arc Set (MWFAS);
while there is a cycle in the graph do
       Identify a cycle using depth-first search;
       Compute the minimum weight among edges in the cycle;
       foreach edge $(u,v)$ in the cycle do
             Reduce its weight by the minimum weight;
             if its weight becomes zero then
                   Remove the edge from the graph;
Sort removed edges based on their weight in decreasing order;
foreach edge $(u,v)$ removed in Phase 1 do
       if re-adding $(u,v)$ does not create a cycle then
             Re-add $(u,v)$ to the graph;
Step 3: Compute Vertex Rankings;
Perform a topological sort of the modified graph;
Assign rankings to vertices based on their positions in the topological order;
Break ties using (the sum of outgoing edge weights- the sum of incoming edges weights)/(total number of incoming and outgoing edges) ;
return Final rankings and metrics (e.g., removed edges and weights);
Algorithm 1 Rank Elements Based on Pairwise Comparisons
The provided algorithm is implemented in a straightforward manner to efficiently solve the ranking problem using the Minimum Weighted Feedback Arc Set (MWFAS) approach. The graph is initially represented using an adjacency list and a weights dictionary, which allows efficient traversal and edge weight management. The process starts by reading the input graph, then cycles are iteratively removed using a depth-first search (DFS) approach to ensure the graph becomes a Directed Acyclic Graph (DAG). We use DFS to effectively check and re-add edges, maintaining connectivity without creating cycles. The use of a heap-based topological sort helps generate a ranking for the vertices, while any ties are broken using the sum of outgoing edge weights. By combining DFS and heap structures, this approach maintains both simplicity and efficiency, offering a practical solution to handle large graphs while minimizing computational overhead.
3 Experimental Results
The tables below show the results of the experiments. The datasets are some of the datasets used in [18]. The experiments are done on a Macbook Pro with processor 3.3 GHz Dual-Core Intel Core i7 and operating system macOS Monterey version 12.7.6. The code is implemented in Python 3.12.4.
Dataset Name Number of Vertices Number of Edges
England_2009_2010 20 164
England_2010_2011 20 161
England_2011_2012 20 165
England_2012_2013 20 153
England_2013_2014 20 165
England_2014_2015 20 107
Average 20 152.5
Table 1: Number of vertices and edges for each Football dataset.
Dataset Name Naive Loss Simple Loss Ratio Loss Time (s) GNNRank Naive Loss GNNRank Simple Loss GNNRank Ratio Loss
England_2009_2010 0.13 0.53 0.63 0.01 0.15 0.61 0.46
England_2010_2011 0.23 0.94 0.87 0.04 0.20 0.95 0.65
England_2011_2012 0.18 0.72 0.75 0.04 0.29 0.80 0.53
England_2012_2013 0.17 0.70 0.75 0.04 0.21 0.80 0.51
England_2013_2014 0.13 0.55 0.65 0.03 0.14 0.56 0.46
England_2014_2015 0.19 0.78 0.78 0.05 0.24 0.96 0.69
Average 0.17 0.70 0.74 0.04 0.21 0.78 0.55
Table 2: Loss function values and running time for each Footbal dataset.
Dataset Name Number of Vertices Number of Edges
Basketball_1985 282 2904
Basketball_1986 283 2937
Basketball_1987 290 3045
Basketball_1988 290 3099
Basketball_1989 293 3162
Basketball_1990 292 3192
Basketball_1991 295 3218
Basketball_1992 298 3238
Basketball_1993 298 3088
Basketball_1994 301 3144
Basketball_1995 302 3182
Basketball_1996 305 3256
Basketball_1997 305 $3$
Basketball_1998 306 3321
Basketball_1999 310 3385
Basketball_2000 318 3475
Basketball_2001 318 3405
Basketball_2002 321 3505
Basketball_2003 327 3560
Basketball_2004 326 3527
Basketball_2005 330 3622
Basketball_2006 334 3695
Basketball_2007 336 3974
Basketball_2008 342 4051
Basketball_2009 347 4155
Basketball_2010 347 4133
Basketball_2011 345 4086
Basketball_2012 345 4126
Basketball_2013 347 4153
Basketball_2014 351 4196
Table 3: Number of vertices and edges for each basketball dataset.
Dataset Name Naive Loss Simple Loss Ratio Loss Time (s) GNNRank Naive Loss GNNRank Simple Loss GNNRank Ratio Loss
Basketball_1985 0.14 0.58 0.55 0.97 0.18 0.71 0.46
Basketball_1986 0.15 0.60 0.53 0.87 0.17 0.69 0.42
Basketball_1987 0.16 0.64 0.58 1.21 0.19 0.77 0.48
Basketball_1988 0.15 0.62 0.56 1.21 0.18 0.70 0.45
Basketball_1989 0.15 0.61 0.56 1.25 0.17 0.70 0.46
Basketball_1990 0.14 0.57 0.55 1.00 0.17 0.70 0.44
Basketball_1991 0.15 0.62 0.56 1.19 0.18 0.70 0.45
Basketball_1992 0.14 0.57 0.54 1.09 0.17 0.67 0.43
Basketball_1993 0.14 0.57 0.55 0.86 0.17 0.68 0.44
Basketball_1994 0.13 0.54 0.52 0.86 0.17 0.69 0.42
Basketball_1995 0.14 0.58 0.55 0.87 0.18 0.72 0.44
Basketball_1996 0.16 0.62 0.58 0.90 0.19 0.77 0.48
Basketball_1997 0.16 0.62 0.58 0.91 0.19 0.76 0.48
Basketball_1998 0.14 0.57 0.57 1.00 0.18 0.74 0.46
Basketball_1999 0.16 0.63 0.58 0.91 0.18 0.74 0.48
Basketball_2000 0.15 0.59 0.59 0.99 0.19 0.77 0.50
Basketball_2001 0.14 0.57 0.58 1.03 0.18 0.71 0.47
Basketball_2002 0.15 0.62 0.60 0.97 0.18 0.73 0.48
Basketball_2003 0.16 0.64 0.59 1.02 0.19 0.78 0.51
Basketball_2004 0.15 0.59 0.57 1.03 0.17 0.69 0.46
Basketball_2005 0.15 0.58 0.59 1.05 0.18 0.73 0.49
Basketball_2006 0.15 0.60 0.60 1.10 0.19 0.74 0.49
Basketball_2007 0.15 0.61 0.60 1.16 0.19 0.77 0.49
Basketball_2008 0.16 0.63 0.62 1.33 0.20 0.78 0.51
Basketball_2009 0.15 0.60 0.59 1.18 0.18 0.72 0.47
Basketball_2010 0.15 0.60 0.59 1.26 0.18 0.71 0.49
Basketball_2011 0.16 0.62 0.60 1.27 0.18 0.74 0.49
Basketball_2012 0.14 0.55 0.59 1.53 0.17 0.69 0.48
Basketball_2013 0.15 0.60 0.61 1.27 0.19 0.76 0.51
Basketball_2014 0.16 0.63 0.60 1.28 0.19 0.75 0.49
Table 4: Loss function values and running time for each basketball dataset.
Dataset Name Number of Vertices Number of Edges Naive Loss Simple Loss Ratio Loss Time (Seconds)
Business FM Full 113 1787 0.09 0.37 0.57 0.21
Business FM Assoc 113 1416 0.06 0.26 0.54 0.05
Business FM Asst 113 1345 0.07 0.28 0.60 0.08
Business FM All 113 2796 0.09 0.39 0.55 0.15
Average 113 1836 0.08 0.32 0.55 0.12
Table 5: Summary of Business FM datasets and results.
Dataset Name Number of Vertices Number of Edges Naive Loss Simple Loss Ratio Loss Time (Seconds)
Business F Full 113 427 0.05 0.23 0.61 0.03
Business F All 113 1196 0.09 0.37 0.58 0.06
Business F Assoc 113 500 0.02 0.10 0.54 0.04
Business F Asst 113 612 0.07 0.30 0.61 0.03
Average 113 683.75 0.06 0.25 0.58 0.04
Table 6: Summary of Business F datasets and results.
Dataset Name Number of Vertices Number of Edges
HeadtoHead 602 5002
Table 7: Number of vertices and edges for the HeadtoHead dataset.
Dataset Name Naive Loss Simple Loss Ratio Loss Time (s) GNNRank Naive Loss GNNRank Simple Loss GNNRank Ratio Loss
HeadtoHead 0.18 0.75 0.75 4.67 0.24 0.96 0.68
Table 8: Loss function values and running time for the HeadtoHead dataset.
Dataset Name Number of Vertices Number of Edges Naive Loss Simple Loss Ratio Loss Time (Seconds)
ComputerScience_F_All 206 579 0.02 0.08 0.59 0.05
ComputerScience_F_Assoc 206 234 0.00 0.01 0.54 0.04
ComputerScience_F_Asst 206 172 0.00 0.00 0.53 0.06
ComputerScience_F_Full 206 223 0.01 0.05 0.54 0.03
Average 206 302 0.00 0.03 0.55 0.05
Table 9: Summary of Computer Science F datasets and results.
Dataset Name Number of Vertices Number of Edges Naive Loss Simple Loss Ratio Loss Time (Seconds)
History_FM_All 145 2204 0.04 0.19 0.52 0.21
History_FM_Assoc 145 1140 0.02 0.09 0.54 0.07
History_FM_Asst 145 681 0.02 0.08 0.50 0.03
History_FM_Full 145 1204 0.04 0.18 0.58 0.07
Average 145 1307.25 0.03 0.13 0.53 0.09
Table 10: Summary of History FM datasets and results.
The results from our algorithm simulation provide an insightful comparison to the methods discussed in [18]. Below, we analyze these results in detail:
Efficiency and Speed: Our results, as presented in the ”Time (Seconds)” column, demonstrate the efficiency of our approach across all datasets. With an average computation time between 0.02 and 0.35 seconds, our algorithm shows significant advantages in terms of speed. Compared to the methods mentioned in [18], such as [1] and [14], which can be computationally expensive, our algorithm provides substantially faster performance even without leveraging deep learning frameworks.
Loss Metrics: The ”Naive Loss” and ”Simple Loss” values indicate the quality of the rankings produced by our algorithm. In [18], differentiable loss functions such as ”upset, naive” and ”upset, ratio” are introduced to minimize ranking upsets. When compared to these metrics, our ”Naive Loss” and ”Simple Loss” values are both competitive and consistently low. For example, the average ”Simple Loss” across datasets often remains below 1.0, suggesting that our heuristic-based approach maintains strong agreement with the pairwise comparisons, while also being much simpler to implement and computationally efficient.
This is the exact definition of naive upset loss, simple upset loss, and ratio upset loss based on [18]:
  • Upset Naive: The naive upset loss, denoted as $L_{upset,naive}$, is defined to measure the fraction of pairwise relationships whose inferred ranking is in the opposite direction to what the original pairwise measurement indicates. Formally, given the definitions:
    • $M_{1}=A-A^{T}$, where $A$ is the pairwise comparison matrix and $M_{1}$ is a skew-symmetric matrix representing the difference between the pairwise comparisons.
    • $T_{1}=r1^{T}-1r^{T}\in\mathbb{R}^{n\times n}$, where $r\in\mathbb{R}^{n}$ represents the ranking scores of the items, and $T_{1}$ represents the difference in ranking scores for each pair of items.
    • $t$ is the number of nonzero elements in $M_{1}$.
    The naive upset loss is given by:
    $L_{upset,naive}=\sum_{i,j:M_{1}(i,j)\neq 0}\left(\text{sign}(T_{1}(i,j))\neq\text{sign}(M_{1}(i,j))\right)/t.$ (1)
  • Upset Simple: The simple upset loss, denoted as $L_{upset,simple}$, extends the naive upset loss by differentiating between ties and opposite signs. Given the sign function acting element-wise on the corresponding entries, the simple upset loss is defined as:
    $L_{upset,simple}=\frac{\lVert\text{sign}(T_{1})-\text{sign}(M_{1})\rVert^{2}_{F}}{t},$ (2)
    where $t$ represents the number of nonzero elements in $M_{1}$. Unlike the naive upset, this metric distinguishes between ties and cases where the predicted signs are opposite.
  • Upset Ratio: The ratio upset loss, denoted as $L_{upset,ratio}$, takes into account not only the sign but also the magnitude of the rankings. It normalizes the differences in ranking scores and pairwise comparison weights to make them comparable. The ratio upset loss is given by:
    $L_{\text{upset, ratio}}=\frac{1}{t}\sum_{i,j:M_{1}(i,j)\neq 0}\left(\frac{T_{1}(i,j)}{T_{2}(i,j)}-\frac{M_{1}(i,j)}{M_{2}(i,j)}\right)^{2},$ (3)
    where:
    • $T_{2}=r1^{T}+1r^{T}+\epsilon\in\mathbb{R}^{n\times n}$ is a normalization term for pairwise score differences, and $\epsilon>0$ is a small positive value used for numerical stability.
    • $M_{2}=A+A^{T}+\epsilon\in\mathbb{R}^{n\times n}$ is a normalization term for pairwise comparison weights.
    • $t$ is the number of nonzero elements in $M_{1}$.
    The ratio upset loss considers both the direction and magnitude of the relationships between nodes and helps to account for more nuanced comparisons that are not purely binary in nature.
We should clarify that in our view, upset naive loss and upset simple loss are not the best functions for qualifying ranking based on weighted comparisons because they do not consider the weight of the violated comparisons; however, we have measured them for our algorithm to compare it with previous works.
Robustness Across Different Datasets: Our approach shows robustness across various types of datasets, including sports (England Football datasets) and academic hiring (Business FM and Computer Science datasets). While [18] emphasizes leveraging trainable similarity matrices to adapt to different graph types, our method, while simpler, consistently achieves low losses and high computational efficiency. This suggests that our heuristic is versatile and performs well irrespective of the dataset characteristics, which aligns with the goal of creating a flexible ranking framework, as described in [18].
Comparative Insights: [18] introduces inductive biases such as the Fiedler vector and proximal gradient steps to improve ranking recovery, especially on more complex datasets. Our approach, although not relying on such advanced machine learning techniques, delivers comparable ranking quality while significantly outperforming these methods in terms of computation time. This highlights an important trade-off between using graph neural networks for adaptive learning versus employing efficient heuristic-based methods that achieve similar or better results in specific cases.
Lets have a look to the execution times for [18] commands. We run some commands that are written on the formal Github repository of [18] in link https://github.com/SherylHYX/GNNRank/blob/main/README.md. Command python ./train.py --all_methods DIGRAC syncRank --dataset animal creates a GNNRank model for animal data using DIGRAC as GNN, also produces results on SyncRank. Executing it took 47 seconds. Command python ./train.py --N 350 --all_methods DIGRAC ib --lr 0.05 creates a GNNRank model for ERO data using both DIGRAC and ib as GNN with 350 nodes, using 0.05 as learning rate. Each epoch in GNNRank has two splits. After 90 seconds, this command had done one split from each of three epochs. The default number of epochs in GNNRank is 1000 epochs. Command python ./train.py --dataset basketball --season 2010 -SP --all_methods baselines_shorter creates a GNNRank model for basketball data in season 2010 using all baselines excluding mvr and also saves predicted results. After 90 seconds, it had done one split from each of seven epochs.
In conclusion, our results suggest that our approach is highly competitive, especially when computational efficiency is prioritized over the adaptability offered by GNN-based methods. Our method’s loss metrics align well with those reported in [18], while providing a notable advantage in processing speed, making it ideal for real-time or large-scale ranking applications.
4 Conclusion and Future Research
In this paper, we explored the relationship between the ranking problem and the Minimum Weighted Feedback Arc Set (MWFAS) problem. We demonstrated how ranking can be effectively modeled as an MWFAS problem, using an approach that balances solution quality and runtime efficiency. Using a heuristic based on the local ratio method, our proposed algorithm was able to provide fast, scalable solutions that maintained competitive accuracy across different datasets.
Our experimental results indicate that our method not only provides a practical solution for ranking large sets of agents, but also outperforms many existing methods in terms of speed. The flexibility to adapt the black-box algorithm used for MWFAS ensures that users can balance the trade-off between performance and computational cost according to their needs.
For future research, we suggest exploring other exact or heuristic methods to solve the MWFAS problem and examining their impact on the resulting ranking quality. Furthermore, investigating improved cycle detection techniques is a possible research topic. Finding alternative tie-breaking mechanisms for ranking of vertices with no directed path between them could further enhance the effectiveness of our ranking method. These directions hold promise for developing even more robust and efficient approaches to solving the ranking problem in diverse and complex settings.
In writing this paper, ChatGPT was used to assist with generating text [24].